Grade 11Algebra

the number of comman tangents to the circle x2+y2-y+=0 and x2+y2+x=0 is

sevit
7 Years agoGrade 11

1 Answer

Arun
7 Years ago
Dear student
 
C1 ( 0, 1/2) ; C2 (-1/2, 0)
 
r1 = 1/2 , r2 = ½
 
C1C2 = 1/sqrt (2)
 
r1 + r2 = 1
 
As r1 + r2 > C1C2
Hence inteesecting circles.
 
Hence no of common tangent = 2